OpenStudy (anonymous):

find the slope and a point on the line of the following equations y=3x-13, y=3x+11, 3y=x+11, 3y=x-13

OpenStudy (anonymous):

match with these options slope=3 point (2,17) slope 1/3 point (19,2) slope 3 point (5,2) slope 1/3 point (17,-2) slope 1/3 point (-5,2)

OpenStudy (jdoe0001):

http://www.algebra-class.com/images/form.gif <---- pretty straightforward notice all given equations are already in slope-intercept form

OpenStudy (jdoe0001):

hint: the y-intercept IS a point on the line

OpenStudy (anonymous):

the options only list one point, how do I find the equation to match that point?

jimthompson5910 (jim_thompson5910):

To see if a point lies on a given line, you plug the coordinates of that point into the equation. Then you determine if the equation is true after simplifying. If the claim is "The point (2,17) lies on the line y=3x-13", then we test that claim like this y=3x-13 17=3*2-13 ... replace x with 2 and y with 17 17 = 6 - 13 17 = -7 ... this equation is false Since 17 = -7 is false, this means (2,17) does NOT lie on the line y = 3x-13

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome